A company with current assets of $100,000 and current liabilities of $50,000 uses $10,000 in cash to pay off a current liability. After this transaction, the company current ratio will equal.
Answer #1

Journal would be:

current liability a/c..Dr $10,000

To Cash $10,000

Hence current liability and cash balance would decrease by $10,000

Hence total current liability=(50,000-10,000)=$40,000

Total Current assets would be=(100,000-10,000)=$90,000

Current ratio=Current assets/current liability

=90,000/40,000

=2.25